OpenStudy (0_0youscareme):

evaluate 3w+(8-v)t if w=4, v=5 and t=2

Vocaloid (vocaloid):

well, start off by replacing the variables (letters) with the matching numbers 3w + (8-v)t = 3(4) + (8-5)2, agreed?

OpenStudy (0_0youscareme):

yes

Vocaloid (vocaloid):

ok, now time for a little set of math rules called, "the order of operations". we use the name PEMDAS to help us remember what order to do our operations in. it stands for: Parentheses Exponents Multiplication/Division Addition/Subtraction

Vocaloid (vocaloid):

so going back to our problem: 3(4) + (8-5)2 we have (8-5) in parentheses, let's simplify that first 8-5 = ?

OpenStudy (0_0youscareme):

3

Vocaloid (vocaloid):

correct, now we move to multiplication 3(4) + 3(2) first we multiply 3(4) and 3(2), which gives us 12 + 6, agreed?

OpenStudy (0_0youscareme):

yup

Vocaloid (vocaloid):

then we finish up with addition

OpenStudy (0_0youscareme):

witch is 18

Vocaloid (vocaloid):

yup, and that's our final answer

OpenStudy (0_0youscareme):

tysm